A rectangular field is 64 m X 25 m. Shawn wants to fence a square field that has the same area as the rectangular field. How long are the sides of the square field?
Answer by checkley79(3341)

The rectangular field has 2*64+2*25=128+50=178 meters.
178/4=44.5 meters is the sides of the square.